If 3x=3y=3z and x+y+z=27√29

Then √x²+y²+z²=???
1
Expert's answer
2020-03-09T14:14:28-0400

Solution. From 3x=3y=3z get x=y=z. As result


"x=y=z=\frac{x+y+z}{3}=9\sqrt{29}"

Find the value x²+y²+z²


"x^2+y^2+z^2=3\times81\times 29"

Therefore


"\sqrt {x^2+y^2+z^2}=9\sqrt{87}"

Answer.


"\sqrt {x^2+y^2+z^2}=9\sqrt{87}"